OpenStudy (anonymous):

if x= 138degrees 36 minutes and 221degrees 24 minutes solve, cos(2x)-sin^2(x/2)+3/4. Been at this for an hour how can express the answer as a fraction?

OpenStudy (anonymous):

Note that 138degrees 36mins and 221 degrees 24mins = x = -(3/4)

ganeshie8 (ganeshie8):

im having hard time interpreting the question :o

ganeshie8 (ganeshie8):

if psble, could u please take snapshot of original question and attach the pic ? :)

OpenStudy (anonymous):

\[\cos(2x)-\sin^2(\frac{ x }{ 2 })+(\frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

\[x=-(\frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

\[ \sin^2\left(\frac x2\right) = \frac{1-\cos(x)}{2} \]

OpenStudy (anonymous):

I still don't know what the point of the question is through.

OpenStudy (anonymous):

I'm just trying to check my work

ganeshie8 (ganeshie8):

wish i could check ur work.. .but im not able to get wat the question is asking us to solve/find :|

OpenStudy (anonymous):

I'll rephrase it

OpenStudy (anonymous):

\[x=-(\frac{ 3 }{ 4 })\] \[\cos(2x)-\sin^2(\frac{ x }{ 2 })+(\frac{ 3 }{ 4 }) = y\] solve for y

OpenStudy (wolf1728):

John the question starts if x= 138degrees 36 minutes and (then something seems missing here) 221degrees 24 minutes

OpenStudy (anonymous):

I meant that x is equal to both 138degrees 36mines and 221degrees 24minutes, but I realized that it's easier if I expressed both of them as -(3/4)

OpenStudy (anonymous):

cos(-(3/4)))

ganeshie8 (ganeshie8):

\(\cos^{-1}(-3/4) = 138^o 36'\)

ganeshie8 (ganeshie8):

got you :)

OpenStudy (anonymous):

I'm trying to substitute all the X into (-3/4)

ganeshie8 (ganeshie8):

\(\cos(2x)-\sin^2(\frac{ x }{ 2 })+(\frac{ 3 }{ 4 }) = y \) basically the question is to evaluate y at \(x = \cos^{-1}(-3/4)\)

ganeshie8 (ganeshie8):

step1 : convert all terms into cos(x)

ganeshie8 (ganeshie8):

use below identites : \(\cos(2x) = 2\cos^2x-1\) \(\sin^2(\frac{x}{2}) = \frac{1-\cos x}{2}\)

ganeshie8 (ganeshie8):

once u have expressed it in terms of cos(x), u can replace cos(x) wid -3/4

OpenStudy (anonymous):

following that I get, (-(3/4))-(7/8)+(3/4)

ganeshie8 (ganeshie8):

okie lemme check.. .

OpenStudy (anonymous):

Thanks I got it

ganeshie8 (ganeshie8):

hmm im getting 0

ganeshie8 (ganeshie8):

\(\large \cos(2x)-\sin^2(\frac{ x }{ 2 })+(\frac{ 3 }{ 4 }) = y\) \(\large 2\cos^2x - 1 -\frac{1-\cos x}{2}+(\frac{ 3 }{ 4 }) = y\) \(\large 2(\frac{-3}{4})^2 - 1 -\frac{1-\frac{-3}{4}}{2}+(\frac{ 3 }{ 4 }) = y\) \(\large \frac{9}{8} - 1 -\frac{7}{8}+(\frac{ 3 }{ 4 }) = y\) ...